# Statistical machine learning tools for probabilistic closures of turbulence models

### v2

@@ 2 Shell models and closure problems
Figure 1: PDFs for $m=8,\ldots,14$ of $|\mathcal{U}_{0}|$ and $|\mathcal{U}_{1}|$
@@ 3.2 Sparse Identification for Non-linear Dynamics
where $\sigma$ is a diffusion coefficient and $W$ is a standard Brownian motion 46. A similar process was conducted in 47. This allows us to write at least part of the closure variables as a stochastic process that solves equation (16) in the Itô sense 46. We then will solve for the closure variable using a simple Euler-Maruyama scheme, and consider two cases: the first, where closure variables are one realization of said stochastic process, and the second, where the closure variables are a mean path of several realizations.
@@ 4 Results
[…]
@@
All simulations which are run until $\tau=50000$, sufficient to collect robust statistics, and with a timestep $d\tau=10^{-3}$, using a second-order Adams-Bashforth scheme.
@@ 4.1 VAE
The VAE trained for this problem, which will be referred from now on as VAE-M, has a total of five layers and learns data of the form $(\log_{2}|\mathcal{U}_{0}|,\log_{2}|\mathcal{U}_{1}|)$. The encoder has two layers, with 64 and 512 neurons, the latent space is two-dimensional with a layer of 128 neurons, and the decoder is a mirror of the encoder, with two layers of 512 and 64 neurons. Activation functions are all ReLU 48 and weights and biases are initialized uniformly as per 49. It was trained in four sessions of $50000$ epochs with batch sizes varying from $32$ to $256$ with an Adam optimizer 50. We can evaluate how well this density estimation is being performed by comparing original training data with data generated by the VAE-M, as can be seen in figure 2.
(a) Component PDF of $|\mathcal{U}_{0}|$
(b) Component PDF of $|\mathcal{U}_{1}|$
(c) Joint PDF of original data
(d) Joint PDF of data generate by the VAE-M
@@
Figure 5: Cutoff is at $s=9$. Performance of closures based on trained VAE-M, which only includes modules, and VAE-P, which also includes phases. Comparison is made based on Sabra statistics of the fully resolved model. Figures 5(a)5(a) show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 5(f)5(h) show energy flux PDFs for the last shells of the reduced model. Figures 5(d) and 5(e) display moments of orders 2 up to 6 for reduced model simulations performed using VAE-M and VAE-P respectively. Both are compared to Sabra moments in solid lines.
@@ 4.2 SINDy
so the second term on the right-hand side of equations 26 and 27 can be computed from the pre-existing data, while the first ones, containing derivatives of moduli, are to be estimated via SINDy.
@@
for $y_{0}(\tau)=|\mathcal{U}_{0}|$ and $y_{1}(\tau)=|\mathcal{U}_{1}|$. The solution for equations 28 and 29 is deterministic. So, for $\sigma=0.05$ we write
@@
where $Y_{0}$ and $Y_{1}$ are one realization of the stochastic process that solves the stochastic differential equations equations 30 and 31. This is now a probabilist model that was not trained to reproduce the statistics presented in figure 1(a), but instead attempts to evolve our closure variables in an approximate dynamics. We will refer to this closure as One-Path.
@@
where $\overline{Y_{0}}=\langle Y_{0}^{j}\rangle_{*}$ and $\overline{Y_{1}}=\langle Y_{1}^{j}\rangle_{*}$ with $j=1,\cdots,10$. Here, $\langle\cdot\rangle_{*}$ is an ensemble average and each pair $\big{(}Y_{0}^{j},Y_{1}^{j}\big{)}$ is one realization of the stochastic process that solves equations 30 and 31, all evolved from the same initial condition until the same final time. This gives us a type of mean path, which is then used as moduli of closure variables to evolve our reduced models in time. This closure is referred to here as Mean-Path. Numerical solutions to the system described in equations 30 and 31 can be seen in figure 6, while the performance of these closures can be seen in figure 7.
Figure 6: Top row shows one realization of the stochastic process that solves equations 30 and 31 in red, with a comparison between the rescaled Sabra variables in black. Time evolution, with an inset for detail, can be seen in figures figures 6(a) and 6(b), while figure 6(c) shows a phase portrait of long time behavior until $\tau=20000$. Figures 6(d) and 6(e) in the bottom row show 15 realizations of the same process in pale red, with the mean path described in equations 34 and 35 displayed in bold red, against the black rescaled Sabra variables, until $\tau=5000$.
@@
Figure 7: Cut-off shell is $s=12$. Performance of closures based on the stochastic version of the approximate dynamics found by SINDy, both using a single realization and taking a mean path. Comparison is made based on Sabra statistics of the fully resolved model. Figures 7(a)7(c) show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 7(f)7(h) show energy flux PDFs for the last shells of the reduced model. Figures 7(d) and 7(e) display moments of orders 2 up to 6 for reduced model simulations performed using VAE-M and VAE-P respectively. Both are compared to Sabra moments in solid lines.
@@ 5 Discussion
The VAEs, in general, show an improvement from previous work using the Hidden Symmetry framework, especially when we note that the PDFs of real parts displayed a gap that was not closed by any of the time-correlated closures written using Gaussian Mixtures Models, including the multi-time models reported in 52. This is an important indicator that the robustness and suitability of the machine learning tool, employed either in the density estimation or in the dynamics recovery, does play an important role in the performance of the closures, as pointed in previous works 29.
From figures 4 and 5 we see that the VAE-based closures performed consistently without the need to be retrained when we change the cut-off shells, which is a feature inherited from the universality of statistics that we achieved due to the change of variables we applied earlier. It is important to note that, when we fix phases at a strictly dissipating value, the structure functions in both cut-offs tend to stay below the Sabra ones, meaning such closures tend to slightly over-dissipate energy, which is visibly improved when we include phases in the modeling. This is also reflected on the PDFs for energy flux, where we can see that the inclusion of phases adds a significant contribution on the left leg of the PDF, which not only agrees more with the baseline Sabra statistics, but the closure that only includes modules can not recover.
@@
In general, what we see in these results is a confirmation of the projection that, on the Hidden-Symmetry framework, more powerful machine learning tools will improve on the accuracy of statistics recovery. It is also remarkable that our closures work for different cutoff shells without the need to retrain any of the machine learning tools used in this work.
However, despite being an improvement on previous work on the same framework, our results can still be improved by, for example, including phases on the SINDy approach, which we did not succeed in doing due to the periodic nature of phases, which reflects as jump-like discontinuities in their time evolution. Our results could also be majorly improved by including time conditioning, either in an alternative type of variational neural network, or by training a conditional VAE.

### v3

@@ 2 Shell models and closure problems
Figure 1: PDFs of $|\mathcal{U}_{0}|$ and $|\mathcal{U}_{1}|$ in 1(a) and of multipliers phases in 1(b) for $m=8,\ldots,14$.
@@ 3.2 Sparse Identification for Non-linear Dynamics
where $\sigma$ is a diffusion coefficient and $W$ is a standard Brownian motion 46. A similar process was conducted in 47. This allows us to write at least part of the closure variables as a stochastic process that solves equation (16) in the Itô sense 46. We then will solve for the closure variable using a simple Euler-Maruyama scheme48, and consider two cases: the first, where closure variables are one realization of said stochastic process, and the second, where the closure variables are a mean path of several realizations.
@@ 4 Results
for $p=2,\dots,6$, where $\langle\cdot\rangle$ is a time average, as well as energy flux through shell $n$, given by
@@
The final comparison we are reporting is local slopes of $p$-order moments, given by
$\zeta_{n}^{p}=\frac{\log(S_{p}(k_{n+1}))-\log(S_{p}(k_{n}))}{\log(\lambda)}.$ (20)
This quantity is useful to evaluate accuracy of closures based on how they compare against the same quantity $\zeta_{n,S}^{p}$ computed for the Sabra moments of order $p$. Such comparison can be commonly found in the literature 32, 49, 31, here we show local slopes for $p=2$, and normalized by the reference value of $\zeta_{n,S}^{p}$.
Simulations of reduced models using the VAE closures run until $\tau=50000$, sufficient to collect robust statistics, and with a timestep $d\tau=10^{-3}$, using a second-order Adams-Bashforth scheme. Simulations of reduced models using the SINDy closures run for the same length of time with the same timestep, but with an Euler-Maruyama scheme.
In general, to simulate equation (16) for a given approximate dynamic and a chosen $\sigma$, a more careful choice of timestep may be required if $\sigma$ introduces intricacies such as very elaborate time and/or scale dependence. It is possible that these will have an effect on the choice of timestep, or even on the choice of numerical method. Our choice of timestep is sustained by the fact that our $\sigma$ will be chosen as constant in the next section.
@@ 4.1 VAE
(a) Component PDF of $|\mathcal{U}_{0}|$
(b) Component PDF of $|\mathcal{U}_{1}|$
(c) Joint PDF of original data
(d) Joint PDF of data generate by the VAE-M
@@
The VAE trained for this problem, which will be referred from now on as VAE-M, has a total of five layers and learns data of the form $(\log_{2}|\mathcal{U}_{0}|,\log_{2}|\mathcal{U}_{1}|)$. The encoder has two layers, with 64 and 512 neurons, the latent space is two-dimensional with a layer of 128 neurons, and the decoder is a mirror of the encoder, with two layers of 512 and 64 neurons. Activation functions are all ReLU 50 and weights and biases are initialized uniformly as per 51. It was trained in four sessions of $50000$ epochs with batch sizes varying from $32$ to $256$ with an Adam optimizer 52. We can evaluate how well this density estimation is being performed by comparing original training data with data generated by the VAE-M, as can be seen in figure 2.
@@
Figure 5: Cutoff is at $s=9$. Performance of closures based on trained VAE-M, which only includes modules, and VAE-P, which also includes phases. Comparison is made based on Sabra statistics of the fully resolved model. Figures $5(a)$ show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 5(f)5(h) show energy flux PDFs for the last shells of the reduced model. Figures 5(d) and 5(e) display moments of orders 2 up to 6 for reduced model simulations performed using VAE-M and VAE-P respectively. Both are compared to Sabra moments in solid lines.
In figures 4(a)4(c) we see show normalized PDFs of real parts for the cutoff shell $s=12$ and the subsequent closure variables. In figures 4(f)4(h) we see energy flux PDFs for the last shells of the reduced model, computed according to equation (19). In figures 4(d) and 4(e) we see moments of orders 2 up to 6, computed according to equation (18). Note that, as scales progress in the inertial range towards the dissipative regime, Sabra PDFs of real parts become more and more heavy tailed 21, which is also recovered by our model. In energy flux PDFs we can see a clear difference between closures that include phase modeling and closures that only model modules. If no phase modeling is present on a closure, meaning, if multipliers phases are fixed on their most probable (strictly dissipating) value of $\Delta_{0}=\Delta_{1}=\pi/2$, it severely underestimates the back-scattering energy flux. This can be clearly seen in 4(h) by the complete absence of a left leg at the PDF. On the other hand, our closure that includes phases seems to always overcompensate on this matter.
Now, with a change in cutoff, we see in figures 5(a)5(c) normalized PDFs of real parts for the cutoff shell $s=9$ and the subsequent closure variables. In figures 5(f)5(h) we see energy flux PDFs for the last shells of the reduced model as well as moments of orders 2 up to 6 in figures 5(d) and 5(e). It is crucial to note that the change of cutoff was done solely in the simulation of the reduced model. Both VAEs were not retrained and are exactly the same networks that generated the closure variables in the case of cutoff $s=12$. Firstly, we can see that the change of cutoffs does not impact qualitatively the behavior of these closures. We also see, more pronouncedly in figures 4(d) and 4(e), that not including phases in the modeling also underestimates moments, with a slight difference in slope. On the other hand, including phases in the modeling leads to a more accurate estimation of moments, albeit a slightly more oscillating one.
@@ 4.2 SINDy
so the second term on the right-hand side of equations˜27 and 28 can be computed from the pre-existing data, while the first ones, containing derivatives of moduli, are to be estimated via SINDy.
@@
for $y_{0}(\tau)=|\mathcal{U}_{0}|$ and $y_{1}(\tau)=|\mathcal{U}_{1}|$. The solution for equations˜29 and 30 is deterministic. So, for $\sigma=0.05$ we write
@@
where $Y_{0}$ and $Y_{1}$ are one realization of the stochastic process that solves the stochastic differential equations equations˜31 and 32. This is now a probabilist model that was not trained to reproduce the statistics presented in figure 1(a), but instead attempts to evolve our closure variables in an approximate dynamics. We will refer to this closure as One-Path.
@@
where $\overline{Y_{0}}=\langle Y_{0}^{j}\rangle_{*}$ and $\overline{Y_{1}}=\langle Y_{1}^{j}\rangle_{*}$ with $j=1,\cdots,10$. Here, $\langle\cdot\rangle_{*}$ is an ensemble average and each pair $\big{(}Y_{0}^{j},Y_{1}^{j}\big{)}$ is one realization of the stochastic process that solves equations˜31 and 32, all evolved from the same initial condition until the same final time. This gives us a type of mean path, which is then used as moduli of closure variables to evolve our reduced models in time. This closure is referred to here as Mean-Path. Numerical solutions to the system described in equations˜31 and 32 can be seen in figure 6, while the performance of these closures can be seen in figures 7 and 7.
Figure 6: Top row shows one realization of the stochastic process that solves equations˜31 and 32 in red, with a comparison between the rescaled Sabra variables in black. Time evolution, with an inset for detail, can be seen in figures figures˜6(a) and 6(b), while figure 6(c) shows a phase portrait of long time behavior until $\tau=20000$. Figures 6(d) and 6(e) in the bottom row show two realization of the same process in pale red, with the mean path described in equations˜35 and 36 displayed in bold red (ensemble average over 100 realizations), against the black rescaled Sabra variables, until $\tau=5000$.
@@
Figure 7: Cut-off shell is $s=12$. Performance of closures based on the stochastic version of the approximate dynamics found by SINDy, both using a single realization and taking a mean path. Comparison is made based on Sabra statistics of the fully resolved model. Figures 7(a)7(c) show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 7(f)7(h) show energy flux PDFs for the last shells of the reduced model. Figures 7(d) and 7(e) display moments of orders 2 up to 6 for reduced model simulations performed using one realization and a mean of several realizations, respectively. Both are compared to Sabra moments in solid lines.
(a) $n=9$
(b) $n=10$
(c) $n=11$
(d) One Path
(e) Mean path
(f) $n=7$
(g) $n=8$
(h) $n=9$
Figure 8: Cut-off shell is $s=9$. Performance of closures based on the stochastic version of the approximate dynamics found by SINDy, both using a single realization and taking a mean path. Comparison is made based on Sabra statistics of the fully resolved model. Figures 8(a)8(c) show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 8(f)8(h) show energy flux PDFs for the last shells of the reduced model. Figures 8(d) and 8(e) display moments of orders 2 up to 6 for reduced model simulations performed using one realization and a mean of several realizations respectively. Both are compared to Sabra moments in solid lines.
In figures 7(a)7(c) we display normalized PDFs of real parts for the cutoff shell $s=12$ and for the closure variables. We can see that, unlike figures 4(a)4(c), the closure PDFs do not display the tendency of becoming more and more heavy tailed. This is expected, given that these closures were built on data solely from the closure variables, and the approximate dynamic for their evolution is run completely decoupled from the reduced system. It is quite remarkable that, despite not including phases in its modeling, which is clear from the lack of back scattering in the energy flux PDF in figure 7(h), both 7(f) and 7(g) show much better agreement with the left legs of the PDFs than the VAE closures reported in figure 4. In figures 7(d) and 7(e) we can see moments of orders 2 up to 6.
Figures 8(a)8(c) present normalized PDFs of real parts for the cutoff shell $s=9$ and the subsequent closure variables. In figures 8(f)8(h) we see energy flux PDFs for the last shells of the reduced model. Again, we point to the fact that this change in cutoff was performed only for the evolution of reduced model. The closure variables were a different realization of the stochastic process that solves equations˜31 and 32 for the One Path closure and the mean of $100$ different realizations for the Mean path closure. We see that qualitative behavior of these closures remained the same, and we remark that the oscillations we see in figure 7(d) and 7(e) are also present in 8(d) and 8(e). Moreover, such oscillations seem to be at the same distance from the cutoff in both cases which indicates that these small discrepancies are due to proximity to the cutoff.
Now, we look at the performance of two closures, the One Path and the VAE-M, but for a cutoff $s=15$. In figures 9(a)9(c) we see normalized PDFs of real parts for the cutoff shell $s=15$ and the subsequent closure variables. We point that, at $n=18$, for our choices of parameters, Sabra is already approaching dissipation range. It is noteworthy that the VAE-M closure managed to recover quite well a statistic so deep in the inertial range as in figure 9(c). In figures 9(f)9(h) we see energy flux PDFs for the last shells of the reduced model. In figures 9(d) and 9(e) the One Path closure and the VAE-M retained the same close-to-cutoff behavior of presenting a small discrepancy, always at the same distance from the cutoff. Again, we see that the qualitative behavior of these closures remained the same as in cutoffs $s=9$ and $s=12$. It is important to recall that all machine learning tools involved in these results were trained for $s=12$, which means that our reduced models display consistent behavior for scales both larger (smaller $s$) and smaller (larger cutoff $s$) than the one where they were trained.
(a) $n=15$
(b) $n=16$
(c) $n=17$
(d) One Path
(e) VAE-M
(f) $n=13$
(g) $n=14$
(h) $n=15$
Figure 9: Cut-off shell is $s=15$. Performance of closures based on the stochastic version of the approximate dynamics found by SINDy using a single realization and based on the trained VAE-M. Comparison is made based on Sabra statistics of the fully resolved model. Figures 7(a)7(c) show normalized PDFs of real parts for the last shell of the reduced model and for the closure variables. Figures 8(h)7(h) show energy flux PDFs for the last shells of the reduced model. Figures 7(d) and 7(e) display moments of orders 2 up to 6 for reduced model simulations performed using a single realization on the SINDy approximation and VAE-M respectively. Both are compared to Sabra moments in solid lines.
Lastly, in figure 10 we present the local slopes of second order moments, defined in equation (20). We see that there is substantial oscillations around the reference value, which is expected from the general behavior of our models in moments recovery. We can also see some other instances of close-to-cutoff behavior discrepancies.
(a) $s=9$
(b) $s=12$
(c) $s=15$
Figure 10: Local slopes for $p=2$ and different values of cutoff $s$, all normalized by the reference value $\zeta_{n,S}^{p}$.
@@ 5 Discussion
The VAEs, in general, show an improvement from previous work using the Hidden Symmetry framework, especially when we note that the PDFs of real parts displayed a gap that was not closed by any of the time-correlated closures written using Gaussian Mixtures Models, including the multi-time models reported in 54. This is an important indicator that the robustness and suitability of the machine learning tool, employed either in the density estimation or in the dynamics recovery, does play an important role in the performance of the closures, as pointed in previous works 29.
From figure 4 paired with 5, as well as figure 7 paired with 8, we see that the all closures performed consistently without the need to be retrained when we change the cut-off shells to a larger scale (smaller $s$) than the one where the tools were trained. Figure 9 is an important result that points to the fact that the same machine learning tools can also be used to close the reduced systems at smaller scales (larger $s$). This is a feature inherited from the universality of statistics that we achieved due to the change of variables we applied earlier.
It is important to note that, when we fix phases at a strictly dissipating value one the VAE based closures, the structure functions in all cut-offs tend to stay below the Sabra ones, meaning such closures tend to slightly over-dissipate energy, which is visibly improved when we include phases in the modeling. This is also reflected on the PDFs for energy flux, where we can see that the inclusion of phases adds a significant contribution on the left leg of the PDF, which not only agrees more with the baseline Sabra statistics, but the closure that only includes modules can not recover.
@@
In general, what we see in these results is a confirmation of the projection that, on the Hidden-Symmetry framework, more powerful machine learning tools will improve on the accuracy of statistics recovery. In this sense, other works 32, 49, 31, where the authors train their machine learning tools directly on Sabra statistics, present much more accurate closures than ours, recovering statistics and local slopes with remarkable precision. More recently 49 it was pointed out that such closures may face a loss of scale invariance on the multipliers statistics near the cutoff. This makes the universality of statistics found in the Hidden-Symmetry framework much more attractive as a potential application of more robust machine learning tools to closure problems in shell models. We have seen throughout this report that the closures trained on the Hidden-Symmetry framework are easily generalizable to different cutoffs, and the employment of more suitable tools presented significant improvement from earlier results in the same framework.
However, while our results improve upon previous work within the same framework, further enhancements are possible, for instance, by incorporating phases into the SINDy approach (forcing properties such as physical periodicity). This was not achieved here due to the periodic nature of phases, which introduces jump-like discontinuities in their time evolution, complicating their inclusion. One can also investigate further different choices of $\sigma$ for the diffusion term in the stochastic SINDy approach. Our results could also be majorly improved by including time conditioning, either in an alternative type of variational neural network, or by training a conditional VAE.